# Faster Online Computation of String Net Frequency

### v1

@@ 3.2 Recalling Weiner’s Suffix Tree Construction
In this subsection, we briefly recall how Weiner’s algorithm updates $\mathsf{STree}(S)$ for a string $S$ to $\mathsf{STree}(aS)$ with a new character $a$.
@@
In Weiner’s algorithm, we climb up the path from the leaf for $S$ and finds the first ancestor $v$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined. See also Figure 1 for illustration. There are two cases:
(1) If ${\mathsf{W\_link}}_{a}({v})$ is a soft Weiner link, then let $ax$ be the destination node of this Weiner link. Note that $av$ is a proper prefix of $ax$. We continue climbing up the path until finding the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ with $a$. We move to node $az$ using the hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$. The locus for $av$ is on the edge from $az$ to $ax$. We split this edge at the locus for $av$, and insert the new leaf for $aS$. Since $av$ is a node in the updated suffix tree $\mathsf{STree}(aS)$, the Weiner link ${\mathsf{W\_link}}_{a}({v})$ becomes hard and points to $av$. The soft Weiner links of the nodes between $v$ and $z$ exclusive are redirected to $av$. Finally, we copy all (soft or hard) Weiner links of node $ax$ to the new node $av$, namely, if ${\mathsf{W\_link}}_{d}({ax})=w$ for character $d\in\Sigma$ and node $w=dax$, then we set ${\mathsf{W\_link}}_{d}({av})=w$.
@@ 3.3 Computing $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$
Suppose that we have built $\mathsf{STree}(S)$ for string $S$ of length $n$ augmented with soft and hard Weiner links, and that we have computed the list $\mathsf{NF^{+}}(S)$. For each node $u$ of $\mathsf{STree}(S)$ with $\phi_{S}(u)\geq 1$, each element $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$ is stored in $u$. By Lemma 4, we can store $\mathsf{NF^{+}}(S)$ for all nodes $u$ of $\mathsf{STree}(S)$ in total $O(n)$ space.
@@
By the definition of NF, we obtain the following:
@@
For any characters $a,b\in\Sigma$, $(a,b)\in\Phi_{S}(u)$ iff
@@
    $u$ is an internal node of $\mathsf{STree}(S)$,
@@
    $u$ has a soft Weiner link ${\mathsf{W\_link}}_{a}({u})$ that points to a leaf, and
@@
    $u$ has an out-edge that begins with $b$ and leads to a leaf.
@@
Suppose $(a,b)\in\Phi_{S}(u)$ holds. Then, by definition, (1) $u$ is a repeat in $S$. Also, $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=\mathsf{occ}_{S}(aub)=1$. By Lemma 1 and by our assumption that $S$ terminates with a unique end-marker \$currency-dollar\\$\$, $u$ is an internal node in $\mathsf{STree}(S)$. Thus, (3) $u$ has an out-edge that begins with $b$ and leads to a leaf. It is clear that $u$ has a Weiner link with $a$. Assume for a contrary that this Weiner link is hard. Then $au$ must be a node in $\mathsf{STree}(S)$. Since $\mathsf{occ}_{S}(au)=1$, $au$ has to be a leaf. But this contradicts that $\mathsf{occ}_{S}(aub)=1$. Thus, (2) ${\mathsf{W\_link}}_{a}({u})$ is a soft Weiner link from $u$. Since $\mathsf{occ}_{S}(au)=1$, this Weiner link points to a leaf.
Suppose (1), (2), and (3) all hold. Then, it is clear that $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$ and $\mathsf{occ}_{S}(u)\geq 2$. Assume for a contrary that $\mathsf{occ}_{S}(aub)=0$. Then, since $\mathsf{occ}_{S}(au)=1$, $au$ has to be a leaf. This contradicts that $u$ has a soft Weiner link ${\mathsf{W\_link}}_{a}({u})=aux$ with some non-empty string $x\in\Sigma^{+}$. Hence $\mathsf{occ}_{S}(aub)>0$. Since $\mathsf{occ}_{S}(au)=1$, we have $\mathsf{occ}_{S}(aub)=1$. Therefore, $(a,b)\in\Phi_{S}(u)$. ∎
By the definition of NF, it suffices for us to check the following types of nodes (recall Section 2.3):
  1. (i)
    Existing maximal repeat $u$ in $aS$ that obtains a new left-extension $a$ with $\mathsf{occ}_{aS}(aub)=\mathsf{occ}_{aS}(au)=1$, and has an existing right-extension $b$ with $\mathsf{occ}_{aS}(ub)=1$.
  2. (ii)
    New maximal repeat $w$ in $aS$ that has left-extension $d$ and right-extension $c$ with $\mathsf{occ}_{aS}(dwc)=\mathsf{occ}_{aS}(dw)=\mathsf{occ}_{aS}(wc)=1$.
  3. (iii)-1
    Existing maximal repeat $v$ in $aS$ that has a left-extension $a$ with $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$.
  4. (iii)-2
    Existing maximal repeat $y$ in $aS$ that has a right-extension $e$ with $\mathsf{occ}_{S}(ye)=1$ and $\mathsf{occ}_{aS}(ye)>1$.
@@
3.4 Dealing with Type (i) node
Here we show that the number of nodes in Type (i) is at most one per iteration. See also Figure 1 for an illustration.
@@ 3.4 Dealing with Type (i) node
If $(u,\phi_{aS}(u))$ is a new element in $\mathsf{NF^{+}}(aS)$ for an existing node $u$, then $u$ is the parent of the leaf representing $S$.
@@
Assume for a contrary that $(u^{\prime},\phi_{aS}(u^{\prime}))\in\mathsf{NF^{+}}(aS)\setminus\mathsf{NF^{+}}(S)$ for an existing node $u^{\prime}$ that is not the parent of the leaf for $S$, namely $u^{\prime}\neq u$. By assumption, $\mathsf{occ}_{aS}(au^{\prime})=1$ and $\mathsf{occ}_{S}(au^{\prime})=0$. This means that $u^{\prime}$ is a node on the path $P$ between $v$ and the leaf for $S$, exclusive. Since $u^{\prime}$ is a non-parent ancestor of the leaf for $S$, we have $\mathsf{occ}_{S}(u^{\prime}b^{\prime})\geq 2$, where $b^{\prime}$ is the first character of the out-edge of $u^{\prime}$ on the path $P$. By Lemma 3, $a$ is the unique paring character for $b^{\prime}$, namely $(a,b^{\prime})\in\Phi_{aS}(u^{\prime})$. However, since $\mathsf{occ}_{S}(u^{\prime}b^{\prime})\geq 2$, we have $(a,b^{\prime})\notin\Phi_{aS}(u^{\prime})$, a contradiction. ∎
3.5 Dealing with Type (ii) node
Since Weiner’s algorithm inserts at most one new internal node per iteration, the number of nodes in Type (ii) is at most one per iteration. Let $av$ be the parent of the new leaf representing $aS$. Then, there exists a unique character $d$ that immediately precedes $ax$ in $S$. The following lemma shows the only case where the node $av$ obtains a new pair of characters in $\Phi_{aS}(av)$. See also Figure 2 for an illustration.

### v2

@@ 3.2 Recalling Weiner’s Suffix Tree Construction
In this subsection, we briefly recall how Weiner’s algorithm updates $\mathsf{STree}(S)$ for a string $S$ to $\mathsf{STree}(aS)$ with a new character $a$ that is prepended to $S$.
@@
In Weiner’s algorithm, we climb up the path from the leaf for $S$ and find the first ancestor $v$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined. See also Figure 1 for illustration. There are two cases:
(1) If ${\mathsf{W\_link}}_{a}({v})$ is a soft Weiner link, then let $ax$ be the destination node of this Weiner link. Note that $av$ is a proper prefix of $ax$. We continue climbing up the path until finding the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ with $a$. We move to node $az$ using the hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$. The locus for $av$ is on the edge from $az$ to $ax$. We split this edge at the locus for $av$, and insert the new leaf for $aS$. Since $av$ is a node in the updated suffix tree $\mathsf{STree}(aS)$, the Weiner link ${\mathsf{W\_link}}_{a}({v})$ becomes hard and points to $av$. The soft Weiner links of the nodes between $v$ and $z$ exclusive are redirected to $av$. Finally, we copy all (soft or hard) Weiner links of node $ax$ to the new node $av$, namely, if ${\mathsf{W\_link}}_{d}({ax})=y$ for character $d\in\Sigma$ and node $y=dax$, then we set ${\mathsf{W\_link}}_{d}({av})=y$.
@@ 3.3 Computing $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$
Suppose that we have built $\mathsf{STree}(S)$ for string $S$ of length $n$ augmented with soft and hard Weiner links, and that we have computed the list $\mathsf{NF^{+}}(S)$. For each node $u$ of $\mathsf{STree}(S)$ with $\phi_{S}(u)\geq 1$, each element $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$ is stored in $u$. By Lemma 1 (or alternatively by Lemma 4), we can store $\mathsf{NF^{+}}(S)$ for all nodes $u$ of $\mathsf{STree}(S)$ in total $O(n)$ space.
@@
By recalling the definitions of $\Phi_{S}(w)$ and $\phi_{S}(w)$ respectively in Equation 1 and in Equation 2 for a string $w$, we can reduce Problem 1 to computing ordered character pairs $(\alpha,\beta)\in\Sigma\times\Sigma$ as described in the following lemma:
@@
For any characters $\alpha,\beta\in\Sigma$ and strings $w,S\in\Sigma^{*}$, $(\alpha,\beta)\in\Phi_{S}(w)$ iff
@@
    $w$ is an internal branching node of $\mathsf{STree}(S)$,
@@
    $w$ has a soft Weiner link ${\mathsf{W\_link}}_{\alpha}({w})$ that points to a leaf $\alpha wt$, where $t\in\Sigma^{+}$ and $\beta=t[1]$, and
@@
    $w$ has an out-edge that begins with $\beta$ and leads to a leaf.
@@
Suppose $(\alpha,\beta)\in\Phi_{S}(w)$ holds. Then, by definition, $w$ is a repeat in $S$. Further, by Lemma 1 and by our assumption that $S$ terminates with a unique end-marker \$currency-dollar\\$\$, (1) $w$ is an internal branching node in $\mathsf{STree}(S)$. By assumption we have $\mathsf{occ}_{S}(w\beta)=1$, and thus (3) $w$ has an out-edge that begins with $\beta$ and leads to a leaf. Also, by assumption we have $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(\alpha w\beta)=1$. This implies that $\alpha w$ is on an edge leading to a leaf $\alpha wt$ with $t\in\Sigma^{+}$ and $t[1]=\beta$. Also, $\alpha w\beta$ is on the same edge if $|t|>1$, or it is a leaf if $|t|=1$. Since $w$ is preceded by $\alpha$ in $S$, $w$ has a Weiner link ${\mathsf{W\_link}}_{\alpha}({w})$ labeled $\alpha$. Since $\alpha w$ is on an edge, ${\mathsf{W\_link}}_{\alpha}({w})=\alpha wt$ with $|t|\geq 1$, and thus it is a soft Weiner link. This leads to (2).
Suppose (1), (2), and (3) all hold. Then, it is clear that $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{S}(w)\geq 2$. ∎
Figure 2: Illustration for Lemma 6, where the rectangles represent leaves, the solid circles represent internal nodes, and the dashed circle represents an implicit node (a locus on an edge). The dashed arc represents a soft Weiner link.
See Figure 2 that illustrates the situation in Lemma 6. The intuition of our online NF computation based on Weiner’s algorithm is as follows: Each time the suffix tree is updated, we examine whether a substructure described in Figure 2 newly appears in the suffix tree, and whether an existing substructure described in Figure 2 disappears from the suffix tree. More formally, by the definitions of $\mathsf{NF^{+}}(S)$ and $\mathsf{NF^{+}}(aS)$ and by Lemma 6, it suffices for us to check the following types of nodes, where Type (i), Type (ii), and Type (iii)-1&2 correspond to Case (i), Case (ii), and Case (iii) of Section 2.3, respectively:
Type (i)
Existing maximal repeat $w$ in $aS$ that obtains a new left-extension $\alpha$ and an existing right-extension $\beta$ such that $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w)=\mathsf{occ}_{aS}(w\beta)=1$. Notice that $w\in\mathsf{M}(S)$ since $\phi_{S}(w)\geq 1$ for Case (i) in Section 2.3. We will later show that the parent $u$ of the leaf for $S$ is the only candidate for this type.
Type (ii)
New maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(w\beta)=1$. We will later show that $w\notin\mathsf{M}(S)$ and $w\in\mathsf{M}(aS)$ for any Type (ii) node $w$, and that the parent $av$ of the new leaf for $aS$ is the only candidate for this type.
Type (iii)-1
Existing maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(\alpha w)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-1, since $\phi_{S}(w)\geq 1$ for Case (iii) in Section 2.3. We will later show that the lowest ancestor $v$ of the leaf for $S$ that has a soft Weiner link labeled $a$ in $\mathsf{STree}(S)$ is the only candidate for this type.
Type (iii)-2
Existing maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(w\beta)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-2, since $\phi_{S}(w)=1$ for Case (iii) in Section 2.3. We will later show that the node $az$ that is the destination node of the hard Weiner link from $z$ labeled $a$ is the only candidate for this type.
In Sections 3.4-3.7, we show how to deal with the nodes of Types (i), (ii), (iii)-1, and (iii)-2. Our algorithm simply uses Weiner’s algorithm that updates $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$, and it suffices for us to perform simple verifications for the visited nodes based on Lemma 6.
3.4 Dealing with Type (i) node
Here we show that the number of nodes of Type (i) is at most one per iteration.
Lemma 7.
Let $u$ be the parent of the leaf representing $S$ in $\mathsf{STree}(S)$. If $u\neq v$, then $u$ is a node of Type (i) and we have $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$, where $b$ is the first character of the edge label from $u$ to $S$ (see also Figure 1). In addition, $u$ is the only possible node of Type (i).
Proof.
It follows from the definition of Type (i) nodes that for any Type (i) nodes, the first character $\alpha$ in the pair $(\alpha,\beta)$ has to be the new character $a$ that is prepended to $S$. If $(a,\beta)\in\Phi_{aS}(u)$ for the parent $u$ of the leaf $S$, then this character $\beta$ must be unique by Lemma 3. Further by Lemma 6, this $\beta$ is the first character $b$ of the edge label from $u$ to $S$ (see also the right diagram of Figure 1). Now consider the case that $u\neq v$, where $v$ is the lowest ancestor of $S$ that has a (soft or hard) Weiner link labeled $a$ in $\mathsf{STree}(S)$ before the update. Since $u\neq v$, $u$ obtains a new soft Weiner link ${\mathsf{W\_link}}_{a}({u})$ labeled $a$ toward the leaf $aS$ in $\mathsf{STree}(aS)$ after the update (see the right diagram of Figure 1). Then the conditions of Lemma 6 are satisfied for the node $u$ and the character pair $(a,b)$ in $\mathsf{STree}(aS)$. Therefore, $(a,b)\in\Phi_{aS}(u)$ holds in this case. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Clearly $(a,b)\notin\Phi_{S}(u)$ because $\mathsf{occ}_{S}(aub)=0$. Hence $u$ is of Type (i) when the aforementioned conditions are satisfied.
Next, we show that $u$ is the only possible node of Type (i). Let $u^{\prime}$ be any node on the path between $v$ and the leaf for $S$, exclusive. Notice that these nodes $u^{\prime}$ are the only nodes other than $u$ that obtain a new left-extension $a$ (see also Figure 1). Assume for a contrary that $u^{\prime}$ is of Type (i), which implies that $\mathsf{occ}_{aS}(au^{\prime})=1$ and $\mathsf{occ}_{S}(au^{\prime})=0$. Let $b^{\prime}$ be the only character that immediately follows $au^{\prime}$ in $aS$, namely $\mathsf{occ}_{aS}(au^{\prime}b^{\prime})=1$. Note that such character $b^{\prime}$ must exist for each $u^{\prime}$. It follows from Lemma 6 that it suffices to check whether $u^{\prime}$ has an out-edge labeled $b^{\prime}$ that leads to a leaf. However, this is not the case: $au^{\prime}b^{\prime}$ is a prefix of $aS$, and thus $u^{\prime}b^{\prime}$ is a prefix of $S$. Thus, $b^{\prime}$ is the first character in the path label from $u^{\prime}$ to the leaf $S$. Since $u^{\prime}$ is a non-parent ancestor of the leaf $S$, the out-edge from $u^{\prime}$ that begins with $b^{\prime}$ does not lead to a leaf. Thus, $u^{\prime}$ cannot be a Type (i) node by Lemma 6. ∎
This node $u$ is found in the process of climbing up the path from the leaf for $S$ toward its lowest ancestor $v$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({v})$ for $a\in\Sigma$. The character $b$ can easily be obtained by using the string depths of $u$ and $S$. When all the conditions in Lemma 7 are satisfied, then $u$ is of Type (i). We then update $\Phi_{aS}(u)\leftarrow\Phi_{S}(u)\cup\{(a,b)\}$ and replace $(u,\phi_{S}(u))$ with $(u,\phi_{S}(u)+1)=(u,\phi_{aS}(u))$ in the list $\mathcal{L}$.
3.5 Dealing with Type (ii) node
We begin with a characterization of a node $w$ of Type (ii):
Lemma 8.
For any node $w$ of Type (ii), $w\notin\mathsf{M}(S)$ and $w\in\mathsf{M}(aS)$ hold.
Proof.
By Case (ii) in Section 2.3, we have $\phi_{aS}(w)\geq 1$. By Lemma 1, $w\in\mathsf{M}(aS)$ holds.
Suppose for a contrary that $w\in\mathsf{M}(S)$. Then, since $\phi_{S}(w)=0$ (by Case (ii) in Section 2.3), for any pair $(\alpha,\beta)\in\Sigma\times\Sigma$ of characters we have $\mathsf{occ}_{S}(\alpha w)>1$ or $\mathsf{occ}_{S}(w\beta)>1$. By monotonicity, $\mathsf{occ}_{aS}(p)\geq\mathsf{occ}_{S}(p)$ holds for any strings $p,S$ and character $a$. Therefore, we have $\mathsf{occ}_{aS}(\alpha w)>1$ and $\mathsf{occ}_{aS}(w\beta)>1$ for any pair $(\alpha,\beta)$ of characters. However, this contradicts that $w$ has a positive NF in $aS$. ∎
Weiner’s algorithm inserts at most one new internal node per iteration, which is the parent $av$ of the new leaf representing $aS$ (see also Figure 3 for an illustration). Thus, $av$ is the only candidate for a node $w$ of Type (ii), implying that the number of nodes of Type (ii) is at most one per iteration. The following lemma shows the only case where the node $av$ obtains a new pair of characters in $\Phi_{aS}(av)$.
@@ 3.5 Dealing with Type (ii) node
Let $v$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({v})$ is defined (see also Figure 3), and let $ax$ be the destination node of the Weiner link ${\mathsf{W\_link}}_{a}({v})$ from $v$ labeled $a\in\Sigma$ in $\mathsf{STree}(S)$ before the update. Let $c$ be the first character of the edge from $av$ to $ax$. If $av$ is on an edge in $\mathsf{STree}(S)$ and the node $ax$ is a leaf, then $(d,c)\in\Phi_{aS}(av)\setminus\Phi_{S}(av)$, where $d\in\Sigma$ is the unique character which immediately precedes $ax$ in $S$.
@@
Recall that $av$ becomes an internal node as the parent of a new leaf for $aS$ in $\mathsf{STree}(aS)$ after the update. Since Weiner’s algorithm inserts at most one internal node and it is $av$ in this case, the string $dav$ remains on an edge in $\mathsf{STree}(aS)$ after the update. Thus, $av$ obtains a soft Weiner link labeled $d\in\Sigma$ toward node $dax$. Since $ax$ is a leaf, $dax$ is also a leaf. For the string $t$ such that $dax=davt$, we have $t[1]=c$. Thus, we have $(d,c)\in\Phi_{aS}(av)$ by Lemma 6. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Due to Lemma 3, $d$ is the unique character pair for $c$ that can possibly be a member of $\Phi_{aS}(av)$. Since $c$ was the only character immediately following $av$ in $S$, $(d,c)$ is the only character pair that could be a member for $\Phi_{aS}(av)$. Since $\mathsf{occ}_{S}(av)=1$, $av$ was not a repeat in $S$ before the update. Thus $\phi_{S}(av)=0$ and hence $(d,c)\notin\Phi_{S}(av)$ holds for the string $S$ before the update.
If $ax$ is not a leaf, then $\mathsf{occ}_{aS}(avc)\geq\mathsf{occ}_{aS}(ax)>1$, implying $(d,c)\notin\Phi_{aS}(av)$. Also, if $av$ is a node in $\mathsf{STree}(S)$, then $av$ was already a maximal repeat in $S$ and thus $av$ is not a Type (ii) node. ∎
The locus for $av$ is found via the Weiner link ${\mathsf{W\_link}}_{a}({v})$ from the node $v$ during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$. We remark that $av$ is on an edge in $\mathsf{STree}(S)$ and the node $ax$ is a leaf, if and only if ${\mathsf{W\_link}}_{a}({v})$ in $\mathsf{STree}(S)$ before the update is a soft Weiner link. We can easily check if $ax$ is a leaf or not. The character $c$ can be obtained by using the string depths of the nodes $av$ and $ax$ (see also Figure 3). The character $d$ can be found easily (e.g. by using the unique Weiner link from leaf $ax$.)
If the conditions in Lemma 9 are satisfied for the nodes $av,ax$ and characters $d,c$, the conditions in Lemma 6 are also satisfied. Then, we update $\Phi_{aS}(av)\leftarrow\{(d,c)\}$, and add $(av,\phi_{aS}(av))=(av,1)$ to the list $\mathcal{L}$.
3.6 Dealing with Type (iii)-1 node
Here we show that the number of nodes of Type (iii)-1 is also at most one per iteration.